66 kilograms of cement is mixed with 22.5 kilograms of water. However, the mixture is supposed to be 40% water in order to dry properly. Find the amount that should be added.
Please help, thanks!

Answer by stanbon(75887)

66 kilograms of cement is mixed with 22.5 kilograms of water. However, the mixture is supposed to be 40% water in order to dry properly. Find the amount that should be added.
----------------
Start: 22.5/66 = 34% water in the cement before additional water added
---------------
Add: x kg of water
----------------
End up with: (x+66) kg of cement containing 40% water.
---------------
EQUATION:
water + water = water
22.5 + x = 0.40(66+x)
22.5 + x = 26.4 + 0.4x
0.6x = 3.9
x = 6.5 kg of water must be added
==================
Cheers,
Stan H.